# Java-Class-Hijack: Software Supply Chain Attack for Java based on Maven Dependency Resolution and Java Classloading

### v1

@@ 4. Ways to Hijack
Java-Class-Hijack consists in crafting a malicious class and then exploit vulnerabilities in Maven’s design, in order to have the malicious class replace a legitimate one. In this section, we describe two strategies that can lead to the malicious class be incorporated in the dependency tree of an application.
@@ 5. Replication
We present two proofs-of-concept: the first is entirely constructed from scratch to effectively illustrate the phases of the class hijacking attack. The second is based on a real, open-source project available on GitHub (Corona-Warn-App) and is used to highlight the potential impact of the attack.
@@ 5.2. Replication in cwa-server
To assess the potential attack impact on a real project, we replicated the attack on the Corona-Warn-App backend services, referred to as cwa-server. Corona-Warn-App is Germany’s Coronavirus tracking application, used during the 2020 pandemic to perform contact tracing using iOS and Android features. The backend services are written using the popular Java framework Spring-Boot and, since Corona-Warn-App is no longer in use, the repository is currently archived. This makes it an ideal base for a proof-of-concept that would also apply to other Java web applications.
@@ 6. Class Hijacking in Gradle ecosystem
Java-Class-Hijack exploits specific design decisions in Maven related to packaging artifacts, classpath resolution, and the default Java class look up algorithm. In this section we discuss the opportunity for the same attack on an application that builds with Gradle, an alternative build system for Java.
@@ 7. Conclusion & Takeaways
  • Projects built with Gradle are also vulnerable to Java-Class-Hijack, even though Gradle’s design decisions reduce the attack surface.
@@ Acknowledgements.
This work was supported by the CHAINS project funded by the Swedish Foundation for Strategic Research (SSF), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) and the IVADO consortium.

### v2

@@ 4. Ways to Hijack
Java-Class-Hijack consists in crafting a malicious class and then exploit Maven and Java design, in order to have the malicious class replace a legitimate one. In this section, we describe two strategies that can lead to the malicious class be incorporated in the dependency tree of an application.
@@ 5. Replication
We present two proofs-of-concept: the first is entirely constructed from scratch to effectively illustrate the phases of the class hijacking attack. The second is based on a real, open-source project available on GitHub repository called Corona-Warn-App (cwa, 2023) and is used to highlight the potential impact of the attack.
All the proof-of-concepts, attack replication and experiments with other languages are available on our GitHub repository (poc, 2024).
@@ 5.2. Replication in cwa-server
To assess the potential attack impact on a real project, we replicated the attack on the Corona-Warn-App backend services, referred to as cwa-server (cwa, 2023). Corona-Warn-App is Germany’s Coronavirus tracking application, used during the 2020 pandemic to perform contact tracing using iOS and Android features. The backend services are written using the popular Java framework Spring-Boot and, since Corona-Warn-App is no longer in use, the repository is currently archived. This makes it an ideal base for a proof-of-concept that would also apply to other Java web applications.
@@ 6. Class Hijacking in Gradle ecosystem
Java-Class-Hijack exploits specific design decisions in Maven related to packaging artifacts, classpath resolution, and the default Java class look up algorithm. In this section, we discuss the opportunity for the same attack on an application that builds with Gradle, an alternative build system for Java.
@@
7. Mitigation
Even though the industry aims to avoid compromised dependencies in the tree at all, limiting the potential impact of any such dependency on the project should also be an important goal for software engineers. This section presents several possible mitigations for this attack.
7.1. Maven Enforcer Plugin
In Java projects built with Maven, developers could use the maven-enforcer-plugin (mav, 2024) with the banDuplicateClasses (ban, 2024) extra rule enabled. This plugin enforces specific rules at build time and adds extra checks. The banDuplicateClasses rule, in particular, fails the build process when a class collision is detected. However, this rule may increase build time and network usage and can result in false positives, as class collisions are sometimes expected and not indicative of an attack. For example, in the most recent versions of hamcrest, its packages have been published in separated artifacts, while before they were deployed into a single one. When the new version is included, a class collision may happen as Maven dependency resolution is based on the artifact id and not on package id.
7.2. Sealed JARs
A project maintainer can enhance the security of dependent applications by creating a sealed JAR (sea, 2024) for the packages defined by the project itself. When an application uses a sealed JAR and the same package name is redefined, a runtime exception is thrown, thereby reducing the impact of Java-Class-Hijack on the application. However, this is not sufficient, as a complete copy of the original dependency could be included in the malicious dependency, bypassing the sealed JAR definition checks. For example, if the victim only calls org.nice.ClassA, the attacker needs to create a complete new malicious dependency with org.nice.ClassA along with all the classes that are called by org.nice.ClassA in the benign package. If org.nice.ClassA calls org.nice.ClassB, and if the attacker does not include ClassB in the dependency, it will result in a RuntimeException.
7.3. Java Modules
With Java 9, Project Jigsaw (jav, 2024) was introduced into the JDK, adding the concept of modularity to the Java ecosystem. Java Modules facilitate the maintenance of large projects and libraries while enhancing the security of the Java SE Platform. A modularized application is immune to Java-Class-Hijack because a package collision results in an automatic compilation failure. Since defining a module for an application only requires adding a module-info.java file in the project root, this method is recommended to increase security and prevent Java-Class-Hijack. To mitigate the attack, is sufficient that the victim project is defined as a module, the dependencies can also be modules but is not required.
In 4 an example of module definition for our abstract PoC is presented, this results in a compilation failure as org.test.NiceClass is defined in both fakelibrary and nicelibrary.
1module victim {
2 uses org.test.NiceClass;
3 requires org.evil.attackerlibrary;
4 requires fakelibrary;
5 requires nicelibrary;
6}
Listing 4: Example of module-info.java to mitigate the attack
@@ 8. Conclusion & Takeaways
  • Java Modules, when correctly defined, block this type of attack.
  • Projects built with other ecosystem are also vulnerable to Java-Class-Hijack.
@@ Acknowledgements.
We thank Hervé Boutemy, who is a PMC member of The Apache Software Foundation, for giving ideas about mitigations. This work was supported by the CHAINS project funded by the Swedish Foundation for Strategic Research (SSF), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) and the IVADO consortium.